Log 18 to the base 4 is 
(1) rational number
(2) an irrational number
(3) a prime number 
(4) not defined 

Grade:11

1 Answers

Arun
5 years ago
Dear Ananya
 
Log 18 to the base 4 is irrational number.
 
as
let it is a rational number
log4 18 = p/q
 
4^(p/q) = 18
 
2^(2p/q) = 18
 
it is never possible hence it is an irrational number